# Efficient Tool Use with Chain-of-Abstraction Reasoning

### 30 Jan 2024 v1

@@ 2 Related Work
There has been a growing interest in augmenting LLMs using external tools. Considerable work has tried to adapt LLMs as tool-using reasoners through in-context learning, demonstrating promising performance improvements in various applications, e.g., math problem solving (Gao et al., 2023; Chen et al., 2022), biomedical question answering (Jin et al., 2023) and self-critiquing (Gou et al., 2023). Nevertheless, guiding LLMs to effectively use tools using in-context demonstrations is challenging, which requires elaborate task-specific prompt engineering and is restricted by the model’s instruction following ability (Jacovi et al., 2023). Noticing the limitations of in-context learning, several works teach LLMs to learn the usage of tools by fine-tuning (Parisi et al., 2022; Schick et al., 2023; Hao et al., 2023b), which more robustly improves LLMs’ performance. However, all above approaches adopt sequential interactions with tools throughout reasoning, slowing the inference speed as a function of the latency of the tool (or API) and the number of API calls that are made.
Some other prior works focus on using LLMs for multi-step reasoning with other modules. In particular, ReAct (Yao et al., 2023b) and FireAct (Chen et al., 2023) integrate LLMs with tools into a closed loop of thought, action and observation steps. This verbose reasoning loop slows down the LLM decoding, and still incorporates tools via sequential interactions, resulting in inefficient inference. Another line of work, PAL (Gao et al., 2023) and Program of Thoughts (Chen et al., 2022) prompt LLMs to generate program-based reasoning and interact with code executors, which however, heavily rely on closed source coding models, i.e., Codex (Chen et al., 2021), and are restricted to procedural arithmetic reasoning. In our work, we aim to design a more general and efficient strategy for LLMs to leverage tools, especially on multi-step reasoning scenarios.
@@
Several previous work research the planning of tool usage in LLMs. Specifically, HuggingGPT (Shen et al., 2023), Chameleon (Lu et al., 2023), OpenAGI (Ge et al., 2023) and MetaTool (Huang et al., 2023) focus on planning the high-level sequence of using multiple tools to address multi-domain mixed tasks. Similarly, LATM (Cai et al., 2023), ML-BENCH (Liu et al., 2023) and Gorilla (Patil et al., 2023) aim at planning program-level integration of multiple APIs for designing scripts of procedural tasks, e.g., a script for training a model described by a GitHub repository. ToolChain* (Zhuang et al., 2023) combines the planning of tool usage with tree-search-based reasoning (Yao et al., 2023a; Hao et al., 2023a), which is especially useful for procedural tasks (Xu et al., 2023; Cobbe et al., 2021). Different from above work, we focus on the planning of general chain-of-thought (Wei et al., 2022) reasoning with awareness of domain specialized tools.
@@ 3 Method
Our method decouples the general reasoning of LLMs from domain-specific knowledge obtained from external tools. Figure 1 shows an overview of our method. In particular, we first fine-tune LLMs to generate reasoning chains with abstract placeholders, e.g., $y1$, $y2$ and $y3$, [^3] We also test placeholders in single-character format, e.g., $x$, $y$ and $z$, which however leads to sub-optimal results. as shown in Figure 1. In the second stage, we reify each reasoning chain by replacing placeholders with domain-specific knowledge obtained from external tools, e.g., calculation results from a calculator, relevant articles retrieved from web search engine, etc. Finally, the question is answered based on the reified reasoning chain.
@@
Note that an intermediate result may appear multiple times in a re-written answer, e.g., the math calculation result $55$ in Figure 2. We prompt LLaMa-70B to replace all occurrences of the same intermediate result with the same placeholder, thereby explicitly connecting the multiple reasoning steps. To ensure that the re-written data is accurate, we use domain-specialized tools to verify the correctness of each CoA reasoning trace. [^5] Detailed implementations of reasoning chain verification are described in Sec. 4.1 and 4.2. Specifically, we use the tools to execute the labeled operations in each CoA, and only keep questions whose CoA can be infilled with valid results by the tools.
@@ 4.1 Mathematical Reasoning
We use an equation solver to perform the arithmetic derivations required in the math domain. Our equation solver first extracts the derivations labeled in the CoA reasoning, e.g., “[$120+35=y$ 1]" and “[$290-y1=y$ 2]", and combines all derivations into a system of equations. Then the system of equations is solved by the SymPy toolkit, [^7] https://www.sympy.org/en/index.html to get the true value of each variable (i.e., the value of the abstract placeholder). Finally, our equation solver returns the reified chain of reasoning by replacing all the variables with their solved true values (including the answer).
@@
Table 2: An example of CoA fine-tuning data construction in Wiki QA domain.
@@ 4.2 Wikipedia QA
Given a question based on Wikipedia knowledge, the model needs to first identify Wikipedia articles as references related to the question, and then reason on key knowledge in the reference articles to answer the question (as shown in the right column of Figure 1). We assume that the specialized knowledge operation in this domain is the retrieval of relevant Wikipedia articles and important named-entities, which are re-written as Wikipedia searching (WikiSearch) and named-entity recognition (NER) [^8] We use NER to extract entities from the article that bridge the former Wikipedia search results to the latter Wikipedia search queries. queries. Table 2 shows an example of a re-written CoA trace for Wiki QA. [^9] We include more prompting examples of Wiki QA answer re-writing in Appendix C.
@@
We use the HotpotQA (Yang et al., 2018) dataset to construct our fine-tuning CoA data in the Wiki QA domain. HotpotQA contains 113K multi-hop QA examples, each labeled with two Wikipedia articles that provide supporting knowledge. Among the 90447 training QA pairs, we identify 72991 as Bridge QA pairs, where an intermediate entity must be identified to link the answer to the question, as shown in Table 2. The remaining 17456 are Comparison QA pairs, where the attributes of two entities are compared, e.g., “Are Randal Kleiser and Kyle Schickner of the same nationality?”. We prompt LLaMa-70B to re-write these training QAs into CoAs with WikiSearch and NER queries, and verify each CoA with our domain tools (described below), by checking whether all the articles returned by the WikiSearch queries match one of the titles in the gold articles. Finally, 8956 Bridge QAs and 5405 Comparison QAs are selected as fine-tuning data. [^10] Compared to mathematical reasoning, generating CoA data for Wiki QA requires more complex tool use combining WikiSearch and NER models, leading to a lower re-writing success rate ($\sim 15.9\%$). For Wiki QA, we note that besides training a LLM to produce CoA data using WikiSearch, we also fine-tune a second LLM to learn to generate the final gold answer based on a correctly reified CoA reasoning trace.
@@ 4.3 Baselines
We apply our CoA reasoning method to both 7B and 70B LLaMa models, and test various model versions including the first version of LLaMa (Touvron et al., 2023a) and more advanced LLaMa-2 and LLaMa-2-Chat (Touvron et al., 2023b). We compare our method to several baselines, including: a) few-shot prompting using 8 randomly sampled QA exemplars from the original (i.e., not re-written) chain-of-thought data (CoT-FSP), b) fine-tuning with original chain-of-thought data (CoT-FT) [^12] Note that in Wiki QA domain, the HotpotQA data used for prompting or fine-tuning baselines is pre-processed to contain both gold Wikipedia articles (serving as chain-of-thought explanations) and the final answer., and c) Toolformer Schick et al. (2023) which fine-tunes LLMs on CCNet (Wenzek et al., 2020) texts augmented with API calls. For evaluation on Wiki QA, we also compared our method with FireAct (Chen et al., 2023), which fine-tunes LLMs on HotpotQA ReAct (Yao et al., 2022) trajectories distilled from GPT-4 (OpenAI, 2023).
@@ 5.1 Mathematical Reasoning
Table 4 shows the evaluation results for the LLaMa-2 and LLaMa-2-Chat models. [^13] We include similar evaluation results for the original LLaMa model (7B) in Appendix B. On the GSM8K and ASDiv datasets, our chain-of-abstraction (CoA) method outperforms the few-shot baseline method CoT-FSP and the regular fine-tuning baseline CoT-FT, demonstrating that our CoA fine-tuning with tool augmentation is more effective in adapting LLMs to multi-step reasoning tasks. Similarly, when evaluated on SVAMP and MAWPS, CoA also consistently outperforms CoT-FSP. Interestingly, for these out-of-distribution datasets, CoT-FT lags further behind CoA, particularly for 7B models, showing that CoA reasoning yields more distributionally robust reasoning performance.
@@
We verify that the robustness of CoA reasoning does not merely benefit from using additional tools, by fine-tuning another LLM (from the same model backbone) to perform the equation solving instead of calling the equation solver, denoted as CoA (no Tool) in Table 4. We find that CoA (no Tool) performs consistently worse than CoA across all datasets, confirming that using specialized tools enables LLM agents to conduct more precise operations, rather than directly solving the same operations. However, we find that CoA (no Tool) still outperforms all baseline methods on zero-shot generalization to SVAMP and MAWPS datasets, implying that chain-of-abstraction reasoning also contributes to better robustness of CoA, perhaps due to better planning of multiple reasoning steps indexed by abstract variables.
@@
Our findings suggest that the benefits of chain-of-abstraction reasoning are most pronounced when problems require long reasoning chains to be solved. Figure 3 shows the stratified performance of three models on GSM8K QA, relative to the number of reasoning steps in the predicted and gold reasoning chains. Compared to the few-shot CoT-FSP, CoA produces reasoning chains that more often match the length of the gold reasoning chains, as reflected by the heat-map statistics (left column) being more aggregated around the diagonal (comparable to CoT-FT). At the same time, we observe that models achieve better QA accuracy when the number of reasoning steps in their generated answers are aligned with the gold references (i.e., the diagonal of heat-maps in right column). All above results show that fine-tuned models are better at learning to produce reasoning chains that match the true reasoning chain for the problem.
Interestingly, we find that CoA, compared to CoT-FT, achieves higher performance especially on questions that require more reasoning steps. In the right column of Figure 3, CoA’s improvement over CoT-FT is more pronounced on questions with more than $3$ steps in the gold reasoning chain (highlighted with red squares). We also present overall accuracy scores on GSM8K subsets according to varying numbers of gold reasoning steps in Table 5, where we confirm this result, indicating that the model trained with CoA has more robust long chain-of-thought reasoning capability, which is enabled from learning to plan using abstractions.
Figure 3: Fine-grained GSM8K evaluation results on LLaMa-2-Chat-7B w.r.t. the number of reasoning steps in the predicted and gold reasoning chain. (Left) The total number of test examples that belong to each stratum. (Right) The corresponding model accuracy (%) for those examples. Non-diagonal cells with fewer than 15 examples are ignored.
Method Gold Reasoning Step
$\leq 2$ $3$ $4$ $5$ $>5$
CoT-FSP 42.9 26.3 18.0 10.9 3.6
CoT-FT 55.5 42.6 25.8 19.0 10.8
CoA 55.8 44.4 32.5 25.3 15.1
+0.3 +1.8 +6.7 +6.3 +4.3
@@
Importantly, we find that the performance benefits of CoA reasoning do not come with increased computational costs. In Figure 4, we show the average time (seconds) that CoA and baseline agents (seeded with LLaMa-2-Chat-7B) needs to answer a question w.r.t. a varying number of gold reasoning steps. Compared to the CoT baselines, CoA requires less time than the few-shot baseline CoT-FSP, whose generation needs to be conditioned on additional examples. However, CoA is slightly less inference-efficient compared to CoT-FT, likely due to the decoding of additional tokens (e.g., “[” and “]”) for the abstract statements.
@@ 5.2 Wiki QA
Table 7 shows our Wiki QA evaluation results using LLaMa-2-Chat models. [^15] We include similar evaluation results on LLaMa-2-7B in Appendix B. Similar to mathematical reasoning, we fine-tune a new version of Toolformer with in-domain chain-of-thought data from HotpotQA, denoted as Toolformer - Wiki. On HotpotQA, CoA achieves higher exact match rates with the gold reference compared to the few-shot or fine-tuning baselines. In particular, CoA outperforms CoT-FSP, CoT-FT, Toolformer and Toolformer - Wiki on the more challenging bridge-type QAs, where two steps of reasoning over Wikipedia knowledge are consecutively entangled, i.e., cannot be performed independently in parallel as in comparison-type QAs. Compared to FireAct fine-tuning, CoA also achieves better performance on both bridge and comparison QAs, without requiring data distilled from closed source GPT-4.
As with mathematical reasoning, CoA agents also perform more efficient inference than Toolformer and FireAct agents when answering the HotpotQA questions. We also find that CoA is more efficient (Time column) than both CoT-FSP and CoT-FT, as CoA does not require few-shot examples as additional inputs and does not need to generate long Wiki articles, which are instead provided by the Wikipedia search engine. Finally, CoA improves over the baseline methods in zero-shot generalization experiments on other Wiki QA datasets, outperforming all baselines on NaturalQuestions and TriviaQA, and matching the best baselines on WebQuestions.
@@ 6 Conclusion
In this work, we propose to decouple the general reasoning ability of LLM agents from executing specialized knowledge via external tools. Our method, chain-of-abstraction (CoA), encourages LLMs to learn the planning of abstract multi-step reasoning, which are more robust to out-of-distribution knowledge shifts. CoA also achieves a more efficient pipeline for tool usage that significantly improves the speed of tool-augmented multi-step reasoning. The simple, yet effective, implementations of our method on two diverse tasks (i.e., mathematical reasoning and open-domain QA) demonstrate its potential for being adapted to new reasoning scenarios.

### 26 Feb 2024 v2

@@ 2 Related Work
There is growing interest in augmenting LLMs using external tools. Considerable work has tried to adapt LLMs as tool-using reasoners through in-context learning, demonstrating promising performance improvements in various applications, e.g., math problem solving (Gao et al., 2023; Chen et al., 2022), biomedical question answering (Jin et al., 2023) and self-critiquing (Gou et al., 2023). Nevertheless, guiding LLMs to effectively use tools using in-context demonstrations is challenging, which requires elaborate task-specific prompt engineering and is restricted by the model’s instruction following ability (Jacovi et al., 2023). Noticing the limitations of in-context learning, several works teach LLMs to learn the usage of tools by fine-tuning (Parisi et al., 2022; Schick et al., 2023; Hao et al., 2023b), which more robustly improves LLMs’ performance. However, all above approaches adopt sequential interactions with tools throughout reasoning, slowing the inference speed as a function of the latency of the tool (or API) and the number of API calls that are made.
Some other prior works focus on using LLMs for multi-step reasoning with other modules. In particular, ReAct (Yao et al., 2023b) and FireAct (Chen et al., 2023) integrate LLMs with tools into a closed loop of thought, action and observation steps. This verbose reasoning loop slows down the LLM decoding, and still incorporates tools via sequential interactions, resulting in inefficient inference. Another line of work, PAL (Gao et al., 2023) and Program of Thoughts (Chen et al., 2022) prompt LLMs to generate program-based reasoning and interact with code executors, which however heavily rely on closed source coding models, i.e., Codex (Chen et al., 2021), and are restricted to procedural arithmetic reasoning. In our work, we aim to design a more general and efficient strategy for LLMs to leverage tools, especially on multi-step reasoning scenarios.
@@
Several previous works research tool usage planning in LLMs. Specifically, HuggingGPT (Shen et al., 2023), Chameleon (Lu et al., 2023), OpenAGI (Ge et al., 2023) and MetaTool (Huang et al., 2023) focus on planning the high-level sequence of using multiple tools to address multi-domain mixed tasks. Similarly, LATM (Cai et al., 2023), ML-BENCH (Liu et al., 2023) and Gorilla (Patil et al., 2023) aim at planning program-level integration of multiple APIs for designing scripts of procedural tasks, e.g., a script for training a model described by a GitHub repository. ToolChain* (Zhuang et al., 2023) combines the planning of tool usage with tree-search-based reasoning (Yao et al., 2023a; Hao et al., 2023a), which is especially useful for procedural tasks (Xu et al., 2023; Cobbe et al., 2021). Different from above work, we focus on the planning of general chain-of-thought (Wei et al., 2022) reasoning with awareness of domain specialized tools.
@@ 3 Method
Our method decouples the general reasoning of LLMs from domain-specific knowledge obtained from external tools. Figure 1 shows an overview of our method. In particular, we first fine-tune LLMs to generate reasoning chains with abstract placeholders, e.g., $y1$, $y2$ and $y3$, [^3] We also test placeholders in single-character format, e.g., $x$, $y$ and $z$, but these led to sub-optimal results. as shown in Figure 1. In the second stage, we reify each reasoning chain by replacing placeholders with domain-specific knowledge obtained from external tools, e.g., calculation results from a calculator, relevant articles retrieved from web search engine, etc. Finally, the question is answered based on the reified reasoning chain.
@@
Note that an intermediate knowledge operation result may appear multiple times in an answer, e.g., in Figure 2, the first equation’s result $55$ is used in the second equation. We prompt LLaMa-70B to replace all occurrences of the same intermediate result with the same placeholder, thereby explicitly connecting the multiple reasoning steps. To ensure that the re-written data is accurate, we use domain-specialized tools to verify the correctness of each CoA reasoning trace. [^5] Detailed implementations of reasoning chain verification are described in Sec. 4.1 and 4.2. Specifically, we use the tools to execute the labeled operations in each CoA, and only keep questions whose CoA can be infilled with valid results by the tools.
@@ 4.1 Mathematical Reasoning
We use an equation solver to perform the arithmetic derivations required in the math domain. Our equation solver first extracts the derivations labeled in the CoA reasoning, e.g., “[$120+35=y$ 1]" and “[$290-y1=y$ 2]", and combines all derivations into a system of equations. Then the system of equations is solved by the SymPy toolkit, [^7] https://www.sympy.org/en/index.html to get the true value of each variable (i.e., the value of the abstract placeholder). Finally, our equation solver returns the reified chain of reasoning by replacing all the variables with their solved true values (including the final answer).
@@
Table 2: Example of CoA fine-tuning data construction in Wiki QA domain.
@@ 4.2 Wikipedia QA
Given a question based on Wikipedia knowledge, the model needs to first identify Wikipedia articles as references related to the question, and then reason on key knowledge in the reference articles to answer the question (as shown in the right column of Figure 1). We assume that the specialized knowledge operation in this domain is the retrieval of relevant Wikipedia articles and important named-entities, which are re-written as Wikipedia searching (WikiSearch) and named-entity recognition (NER) [^8] We use NER to extract entities from the article that bridge the former WikiSearch results to the latter WikiSearch queries. queries. Table 2 shows an example of a re-written CoA trace for Wiki QA. [^9] We include more prompting examples of Wiki QA answer re-writing in Appendix C.
@@
We use the HotpotQA (Yang et al., 2018) dataset to construct our fine-tuning CoA data in the Wiki QA domain. HotpotQA contains 113K multi-hop QA examples, each labeled with two Wikipedia articles that provide supporting knowledge. Among the 90447 training QA pairs, we identify 72991 as Bridge QA pairs, where an intermediate entity must be identified to link the answer to the question, as shown in Table 2. The remaining 17456 are Comparison QA pairs, where the attributes of two entities are compared, e.g., “Are Randal Kleiser and Kyle Schickner of the same nationality?”. We prompt LLaMa-70B to re-write these training QAs into CoAs with WikiSearch and NER queries, and verify each CoA with our domain tools (described below), by checking whether all the articles returned by the WikiSearch queries match one of the titles in the gold articles. Finally, 8956 Bridge QAs and 5405 Comparison QAs are used as fine-tuning data, whose re-written CoAs pass the verification. [^10] Compared to mathematical reasoning, generating CoA data for Wiki QA requires more complex tool use that combines WikiSearch and NER models, leading to a lower re-writing success rate ($\sim 15.9\%$). For Wiki QA, we note that besides training a LLM to produce CoA data using WikiSearch, we also fine-tune a second LLM to learn to generate the final gold answer based on a correctly reified CoA reasoning trace.
@@ 4.3 Baselines
We apply our CoA reasoning method to both 7B and 70B LLaMa models, and test various model versions including the first version of LLaMa (Touvron et al., 2023a) and the more advanced LLaMa-2 and LLaMa-2-Chat (Touvron et al., 2023b). We compare our method to several baselines, including: a) few-shot prompting using 8 randomly sampled QA exemplars from the original (i.e., not re-written) chain-of-thought data (CoT-FSP), b) fine-tuning with original chain-of-thought data (CoT-FT) [^12] Note that in Wiki QA domain, the HotpotQA data used for prompting or fine-tuning baselines is pre-processed to contain both gold Wikipedia articles (serving as chain-of-thought explanations) and the final answer., and c) Toolformer Schick et al. (2023) which fine-tunes LLMs on CCNet (Wenzek et al., 2020) texts augmented with API calls. For evaluation on Wiki QA, we also compared our method with FireAct (Chen et al., 2023), which fine-tunes LLMs on HotpotQA ReAct (Yao et al., 2022) trajectories distilled from GPT-4 (OpenAI, 2023).
@@ 5.1 Mathematical Reasoning
Table 4 shows the evaluation results for the LLaMa-2 and LLaMa-2-Chat models. [^13] We include similar evaluation results for the original LLaMa model (7B) in Appendix B. On the GSM8K and ASDiv datasets, our CoA method outperforms the few-shot baseline CoT-FSP and the regular fine-tuning baseline CoT-FT, demonstrating that CoA fine-tuning with tool augmentation is more effective in adapting LLMs to multi-step reasoning tasks. Similarly, when evaluated on out-of-distribution datasets, SVAMP and MAWPS, CoA also consistently outperforms the baselines. Interestingly, for these out-of-distribution datasets, CoT-FT lags further behind CoA, particularly for 7B models, showing that CoA reasoning yields more distributionally robust reasoning performance.
@@
We verify that the robust generalization performance of our CoA method does not merely benefit from using additional tools, by fine-tuning another LLM to solve the equation (from the same model backbone), rather than calling the equation solver, denoted as CoA (no Tool) in Table 4. We find that CoA (no Tool) performs consistently worse than CoA across all datasets, confirming that using specialized tools enables LLM agents to conduct more precise operations, rather than directly solving the same operations. However, CoA (no Tool) still outperforms all baseline methods on zero-shot generalization to SVAMP and MAWPS datasets, implying that learning abstract reasoning chains also contributes to better robustness of CoA, perhaps due to better planning of multiple reasoning steps indexed by abstract variables.
@@
Our findings suggest that the benefits of chain-of-abstraction reasoning are most pronounced when problems require long reasoning chains to be solved. Figure 3 shows the stratified performance of three models on GSM8K QA, relative to the number of reasoning steps in the predicted and gold reasoning chains. Compared to the few-shot CoT-FSP, CoA produces reasoning chains that more often match the length of the gold reasoning chains, as reflected by the heat-map statistics (left column) being more aggregated around the diagonal (comparable to CoT-FT). At the same time, we observe that models achieve better QA accuracy when the number of reasoning steps in their generated answers are aligned with the gold references (i.e., the diagonal of heat-maps in right column). Above results show that fine-tuned models are better at learning to produce reasoning chains that match the true reasoning chain for the problem.
Interestingly, we find that CoA, compared to CoT-FT, achieves higher performance especially on questions that require more reasoning steps. In the right column of Figure 3, CoA’s improvement over CoT-FT is more pronounced on questions with more than $3$ steps in the gold reasoning chain (highlighted with red squares). This indicates that the model trained with CoA has more robust long chain-of-thought reasoning capability, which is learned from planning with abstractions.
Figure 3: GSM8K evaluation results on LLaMa-2-Chat-7B w.r.t. the number of reasoning steps in the predicted and gold reasoning chain. (Left) The number of test examples that belong to each stratum. (Right) The corresponding model accuracy (%) for those examples. Non-diagonal cells with fewer than 15 examples are ignored.
@@
Importantly, we find that the performance benefits of CoA reasoning do not come with increased computational costs. In Figure 4, we show the average time (seconds) that CoA and baseline agents (seeded with LLaMa-2-Chat-7B) needs to answer a question w.r.t. required gold reasoning steps. Compared to the CoT baselines, CoA requires less time than the few-shot baseline CoT-FSP, whose generation needs to be conditioned on additional examples. However, CoA is slightly less inference-efficient compared to CoT-FT, likely due to the decoding of additional tokens (e.g., “[” and “]”) for the abstract statements.
@@ 5.2 Wiki QA
Table 6 shows our Wiki QA results using LLaMa-2-Chat models. [^15] We include similar evaluation results on LLaMa-2-7B in Appendix B. Similar to mathematical reasoning, we fine-tune a new version of Toolformer with in-domain chain-of-thought data from HotpotQA, denoted as Toolformer - Wiki. On HotpotQA, CoA achieves higher exact match rates with the gold reference compared to the few-shot or fine-tuning baselines. In particular, CoA outperforms all baselines on the more challenging bridge-type QAs, where two steps of reasoning over Wikipedia knowledge are consecutively entangled, i.e., cannot be performed independently in parallel as in comparison-type QAs. Compared to FireAct fine-tuning, CoA also achieves better performance on both bridge and comparison QAs, without requiring data distilled from closed source GPT-4.
As with mathematical reasoning, CoA agents also perform more efficient inference than Toolformer and FireAct agents when answering HotpotQA questions. We also find that CoA is more efficient (Time column) than both CoT-FSP and CoT-FT, as CoA does not require few-shot examples as additional inputs and does not need to generate long Wiki articles, which are instead provided by the search engine. Finally, CoA improves over the baseline methods in zero-shot generalization experiments on other Wiki QA datasets, outperforming all baselines on NaturalQuestions and TriviaQA, and matching the best baselines on WebQuestions.
@@ 6 Conclusion
In this work, we propose to decouple the general reasoning of LLM agents from specialized knowledge obtained via external tools. Our method, chain-of-abstraction (CoA), encourages LLMs to learn the planning of abstract multi-step reasoning, which are more robust to out-of-distribution knowledge shifts. CoA also achieves a more efficient pipeline for tool usage that significantly improves the speed of tool-augmented multi-step reasoning. The simple, yet effective, implementations of our method on two diverse tasks (i.e., math reasoning and open-domain QA) demonstrate its potential for being adapted to new reasoning scenarios.
Limitations
We acknowledge a few limitations in our work. First, datasets used for testing our method cannot have exhaustive coverage of all reasoning scenarios in real world. We instead consider two representative reasoning domains, i.e., mathematical reasoning and Wikipedia QA, and use English as a primary language in our testing. Furthermore, all models in our experiments use greedy decoding to generate inferences, which leaves room for applying more advanced decoding strategies, e.g., self-consistency (Wang et al., 2022) decoding. In Appendix B, we include an initial study of integrating our CoA reasoning with self-consistency decoding, where our method also achieves promising results, motivating further research on this direction. Finally, our method is tested on the setting of fine-tuning the full LLMs, which requires considerable computational resources, while more efficient model training schemes, e.g., LoRA (Hu et al., 2021), can be applied in future work.
Acknowledgements
We thank Beatriz Borges, Gail Weiss, Syrielle Montariol, Li Mi and Zeming Chen for reading and providing comments on drafts of this paper. Antoine Bosselut gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Science Seed Fund, the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.